OpenStudy (anonymous):

Rewrite each expression without exponential expression: 1.) 16^3/4 2.) x^3/5 3.) (4x+2)^3/2 4.) 1/x^2/3

OpenStudy (anonymous):

1)\[16^{3/4}\]right?

OpenStudy (anonymous):

if yes than\[\sqrt[4]{16^3}= \sqrt[4]{4.096}=8\]